OpenStudy (anonymous):

Determine the roots (zeros) of x^4-3x^3+5x^2-27x-36 = 0 i got one of the zeros from the calculator and it was -1 from that i used synthetic division to get my new equation : x^3-4x^2+9x-36. now i dont know how to get the rest of my zeros. i need help!

OpenStudy (anonymous):

hint: factor out x^2 from the first two terms, and 9 from the second two terms.

OpenStudy (jdoe0001):

hmmm have you covered the "rational root test" yet?

OpenStudy (jdoe0001):

or ... are you supposed to get them from their graph?

OpenStudy (anonymous):

no, my test is tomorrow. this is like a review packet and that is the exact question. @jdoe0001 and @ertdre i did i got (x^2+9) (x-4) that would give me one of my other answers being 4, right? i dont know where to go from there.

OpenStudy (jdoe0001):

so... how are you supposed to find the roots then? what are you meant to use to get them?

OpenStudy (jdoe0001):

or what are you expected to use to get them ?

OpenStudy (here_to_help15):

@jdoe0001 have you not unlocked human calculator at least??

OpenStudy (anonymous):

well im supposed to use synthetic division to get one of the zeros and from the new equation i am supposed to get the rest of my zeros

myininaya (myininaya):

How did you get that polynomial when you assumed one of the zeros was x=-1

myininaya (myininaya):

I'm not getting that polynomial (and I'm not getting a remainder 0)

myininaya (myininaya):

which means x=-1 wasn't a zero

myininaya (myininaya):

unless I have done some mistake

OpenStudy (anonymous):

i got it by doing synthetic division. @myininaya

myininaya (myininaya):

-1| 4 -3 5 -27 -36 | -4 7 -12 39 ------------------------- 4 -7 12 -39 | 3 not remainder 0 so x=-1 is not a zero

OpenStudy (anonymous):

you forgot the 1

myininaya (myininaya):

Oh the coefficent of x^4 is 1 not 4

OpenStudy (anonymous):

yes maam the 4 is the degree

myininaya (myininaya):

-1| 1 -3 5 -27 -36 | -1 4 - 9 36 - -------------------- 1 -4 9 -36 |0 okay x=-1 does work

myininaya (myininaya):

yeah i thought there was a 4 in front also

myininaya (myininaya):

Well you already found the other zeros by factoring that resulting polynomial

myininaya (myininaya):

\[x^3-4x^2+9x-36=0 \\x^2(x-4)+9(x-4)=0 \\ (x-4)(x^2+9)=0 \\ x-4=0 \text{ when } x=?_1 \\ x^2+9=0 \text{ when } x=\pm ?_2\]

OpenStudy (anonymous):

i get everything untill the last line. i dont know how you did that

myininaya (myininaya):

how I got x^2+9?

myininaya (myininaya):

I didn't get that You actually did

myininaya (myininaya):

well and then I also got it after I factored but you got x^2+9 as a factor first

myininaya (myininaya):

Do you mean do hot you solve x^2+9=0 for x?

OpenStudy (anonymous):

i mean your answer to x^2+9

myininaya (myininaya):

\[x^2+9=0 \\ x^2=-9 \] Now you just take the square root of both sides

OpenStudy (anonymous):

i am officially confused

OpenStudy (anonymous):

would it be 3i

myininaya (myininaya):

you would have two solutions

myininaya (myininaya):

+ or - 3i

OpenStudy (jdoe0001):

heheh officially =) so you're "unofficially" doing ok then ? =)

OpenStudy (anonymous):

okay i got it thank you! @myininaya & hah nope now i am officially doing okay because i got exactly what i needed. @jdoe0001

myininaya (myininaya):

so the part that was confusing was just solving the x^2+9=0 thing?

myininaya (myininaya):

Or was it why I set x^2+9=0

myininaya (myininaya):

Or was it where x^2+9 came from?

OpenStudy (anonymous):

it was the solving x^2+9=0 part. i thoight that i was supposed to factor it. i had everything exept that part. i was trying to factor but i saw no way to factor so i was confused

myininaya (myininaya):

\[x^4-4x^3+5x^2-27x-36=0 \\ (x+1)(x-4)(x^2+9)=0\]

myininaya (myininaya):

you can factor x^2+9

myininaya (myininaya):

\[(x-3i)(x+3i)\]

myininaya (myininaya):

\[x^2+a^2 =(x-ai)(x+ai)\]

OpenStudy (anonymous):

i never thought about that though.

myininaya (myininaya):

\[(x-ai)(x+ai)=x^2+xai-aix-a^2i^2 \\ =x^2+0-a^2(-1) \\ =x^2-a^2(-1) \\=x^2+a^2\]

myininaya (myininaya):

You don't have to factor x^2+9 to find the zeros of x^2+9 though

OpenStudy (anonymous):

yeah i know that now.

myininaya (myininaya):

\[x^2+a^2=0 \\ x^2=-a^2 \\ \sqrt{x^2}=\sqrt{-a^2} \\ |x|=\sqrt{-a^2} \\ x=\pm \sqrt{-a^2} \\ x= \pm i \sqrt{a^2} \\x=\pm ai\]